OpenStudy (anonymous):

The amount of money in an account with continuously compounded interest is given by the formula A=Pe^rt, where P is the principal, r is the annual interest rate, and t is the time in years. Calculate to the nearest hundredth of a year how long it takes for an amount of money to double if interest is compounded continuously at 7.2%. Round to the nearest tenth. 9.6 yr 1 yr 7.9 yr 0.6 yr

OpenStudy (anonymous):

You want the money in the account to double, so \(A=2P\): \[2P=Pe^{rt}\\ 2=e^{rt}\] You're given \(r=7.2\%=0.072.\) Solve for \(t\): \[2=e^{0.072t}\\ \ln2=\ln\left(e^{0.072t}\right)\\ \ln2=0.072t\ln e\\ \ln2=0.072t\\ t\approx\cdots\]

OpenStudy (anonymous):

wait would you divide somewhere?

OpenStudy (anonymous):

Yes, \[\ln2=0.072t~\Rightarrow~t=\frac{\ln2}{0.072}\approx\cdots\]

OpenStudy (anonymous):

so 9.6?

OpenStudy (anonymous):

Yep

OpenStudy (anonymous):

thank you can you help me with something else?

OpenStudy (anonymous):

Sure, would you mind posting it as another question? In case I can't answer it, someone else may be able to help you more readily.

OpenStudy (anonymous):

sure